### v1

# K theoretic virtual fundamental cycle of global Kuranishi chart

Abstract.
In this paper, we define the virtual fundamental cycle of a global Kuranishi chart, as an element in the (analytic) orbifold K-homology of the base space of the chart, and verify that it defines the same invariants as in [2].
1. Introduction
The motivation of this paper is to define permutation equivariant quantum K-theory [12][6][7] for general Symplectic target spaces. The essential ingredient in the definition is the $K$-theoretical virtual fundamental cycle of moduli spaces $\mathcal{M}$ of J-holomorphic curves. These moduli spaces are characterized using almost complex global Kuranishi charts [1][2].
Definition 1.1.
An almost complex global Kuranishi chart $\mathcal{A}=(\mathcal{Y},\mathcal{E},\sigma)$ of $\mathcal{M}$ consists of an orbi-section $\sigma$ of a complex orbi-bundle $\mathcal{E}$ over an almost complex orbifold $\mathcal{Y}$ with boundary, such that $\sigma^{-1}(0)=\mathcal{M}\subseteq\mathcal{Y}\setminus\partial\mathcal{Y}$.
In this paper, we define $K$-theoretical virtual fundamental cycles as elements in the $K$-homologies of certain spaces, instead of $K$-cohomologies of cotangent spaces as in [1][2].
1.1. The case of manifolds
Let $Y$ be a compact almost complex manifold (possibly with boundary), and $s$ a section of a complex vector bundle $E$ over $Y$ whose zero locus $s^{-1}(0)$ lies in the interior of $Y$. Pick Hermitian metrics on $T^{\ast}Y$ and $E$.
Let $\pi:T^{\ast}Y\to Y$ and
$\begin{array}[]{llll}S_{+}&=\wedge^{even}T^{\ast}Y,&S_{-}&=\wedge^{odd}T^{\ast}Y,\\ E_{+}&=\wedge^{even}E,&E_{-}&=\wedge^{odd}E.\end{array}$
Given a section $\tau:X\to F$ of a Hermitian bundle $F$ over a manifold $X$, we define $\lambda_{\tau}:\wedge^{\ast}F\to\wedge^{\ast}F$ by $\tau(x)\wedge\cdot+i_{\tau(x)}(\cdot)$ over points $x\in X$, where $i_{\tau(x)}$ is the contraction with $\tau(x)$ via the Hermitian metric on $F$. Form $\lambda_{\xi}:\pi^{\ast}S_{\pm}\to\pi^{\ast}S_{\mp}$ and $\lambda_{s}:E_{\pm}\to E_{\mp}$ corresponding to the diagonal map $\xi:T^{\ast}Y\to\pi^{\ast}T^{\ast}Y$ and $s:Y\to E$.
Define
$\begin{array}[]{llll}O_{+}&=(S_{+}\otimes E_{+})&\oplus&(S_{-}\otimes E_{-}),\\ O_{-}&=(S_{-}\otimes E_{+})&\oplus&(S_{+}\otimes E_{-}),\end{array}$
and
$\alpha=\left[\begin{matrix}\lambda_{\xi}\otimes 1&1\otimes\pi^{\ast}\lambda_{s}\\ 1\otimes\pi^{\ast}\lambda_{s}&-\lambda_{\xi}\otimes 1\end{matrix}\right]:\pi^{\ast}O_{+}\to\pi^{\ast}O_{-}.$
Pick a vector bundle $O^{\prime}$ whose direct sum with $O_{+}$ is trivial, and form $(\alpha^{\prime}:\pi^{\ast}O^{\prime}_{+}\to\pi^{\ast}O^{\prime}_{-})=(\alpha\oplus 1:\pi^{\ast}(O_{+}\oplus O^{\prime})\to\pi^{\ast}(O_{-}\oplus O^{\prime}))$. Since $\alpha^{\ast}\alpha$ is multiplication by $|\xi|^{2}+|s(x)|^{2}$, $\alpha$, hence $\alpha^{\prime}$, is a bundle isomorphism outside the compact set $s^{-1}(0)$. Under a suitable trivialization of $O^{\prime}_{\pm}$, $\alpha^{\prime}$ is the identity map near $\partial Y$. We can also modify $\alpha$ so that it is constant along $\{(x,\nu\xi):\nu\in\mathbb{R}_{+}\}$ over the entire $Y$ by [4] page 492-493.
Let $H_{\pm}=L^{2}(Y,O^{\prime}_{\pm})$. Note that $H_{\pm}$ carry the structure of $\mathbb{C}(Y)$-modules via the natural multiplication by continuous $\mathbb{C}$-valued functions on the base, which can be understood as *-representations $\phi_{\pm}:\mathbb{C}(Y)\to L(H_{\pm})$ to the algebras of bounded operators on the Hilbert spaces. Let $T_{s}:H_{+}\to H_{-}$ be the order 00 pseudo-differential operator obtained by continuously extending the quantization of $\alpha^{\prime}$ from compactly supported smooth sections of $O^{\prime}_{\pm}$ to $H_{\pm}$. The datum $(H_{\pm},\phi_{\pm},T_{s})$ defines an element in Kasparov’s K-homology group $K_{0}(Y)$, which we will denote as $\mathcal{O}^{vir}_{s}$ and refer to as the K-theoretic virtual fundamental class.
In particular, if $s$ is transverse to the zero section, so that $M=s^{-1}(0)$ is a stably almost complex submanifold, such extension construction allows us to define the K-theoretic fundamental class $[M]\in K_{0}(\Sigma^{N}M^{+})=K_{0}(M)$.
Theorem 1.
Let $i:M\to Y$ be the inclusion map. Then $i_{*}[M]=\mathcal{O}^{vir}_{s}$.
Given a virtual bundle $V=V_{+}\ominus V_{-}\in K^{0}(Y)$, we define the pairing $\mathcal{O}^{vir}_{s},V\right>$ as the difference $n_{+}-n_{-}$ of Fredholm indices
$\begin{array}[]{ll}n_{+}&=\operatorname{ind}(T_{s}\otimes 1_{+}:H_{+}\otimes_{\mathbb{C}(Y)}\mathbb{C}(V_{+})\to H_{-}\otimes_{\mathbb{C}(Y)}\mathbb{C}(V_{+}));\\ n_{-}&=\operatorname{ind}(T_{s}\otimes 1_{-}:H_{+}\otimes_{\mathbb{C}(Y)}\mathbb{C}(V_{-})\to H_{-}\otimes_{\mathbb{C}(Y)}\mathbb{C}(V_{-})),\end{array}$
where $\mathbb{C}(V)$ is the space of complex-valued continuous sections of the bundle $V$, and has the structure of a finitely-generated $\mathbb{C}(Y)$-module.
Theorem 2.
The pairing $\mathcal{O}^{vir}_{s},V\right>$ equals the index of the virtual bundle
$O\otimes\pi^{\ast}V\in K^{0}(T^{\ast}Y),$
where $O=O_{+}\ominus O_{-}$.
This theorem shows that our definition agrees with that in [2].
1.2. Results
In this paper, we defined a $K$-theoretic virtual fundamental class $\mathcal{O}^{vir}_{\mathcal{Y,E},\sigma}$ of an almost complex global Kuranishi chart $(\mathcal{Y,E},\sigma)$. The $K$-theoretic virtual fundamental class is an element in $K_{0}(|I\mathcal{Y}|)$, the (complex coefficient) K-homology of the coarse space $|I\mathcal{Y}|$ of the inertia orbifold of $\mathcal{Y}$. We also define the pairing of $\mathcal{O}^{vir}_{\mathcal{Y,E},\sigma}$ and elements in the (orbifold) K-group of $\mathcal{Y}$.
Sector-wise, $\mathcal{O}^{vir}_{\mathcal{Y,E},\sigma}$ is constructed from $(|\mathcal{Y}^{(g)}|,str_{(g)}\mathcal{E}|_{\mathcal{Y}^{(g)}},str_{(g)}\circ\sigma|_{\mathcal{Y}^{(g)}})$ as in the no-isotropy case.
The pairing is the cap product composed with a certain map $p_{\ast}:K_{0}(|I\mathcal{Y}|)\to\mathbb{C}$.
  • The cap product of $(T:H_{+}\to H_{-})\in K_{0}(|I\mathcal{Y}|)$ and $(\alpha:V_{+}\to V_{-})\in K^{0}(|I\mathcal{Y}|)$ is
    $\left[\begin{matrix}T\otimes 1&1\otimes C(\alpha^{\ast})\\ 1\otimes C(\alpha)&-T^{\ast}\otimes 1\end{matrix}\right]:\begin{matrix}H_{+}\otimes C(V_{+})&&H_{-}\otimes C(V_{+})\\ \oplus&\to&\oplus\\ H_{-}\otimes C(V_{-})&&H_{+}\otimes C(V_{-})\end{matrix}.$
    Here in the representatives, $H_{\pm}$ are Hilbert spaces and $V_{\pm}$ are Hermitian bundles, $C(V_{\pm})$ are continuous sections of $V_{\pm}$ and $C(\alpha):C(V_{+})\to C(V_{-})$ the induced map.
  • The map $p_{\ast}$ is an extension of the Fredholm index on $K$-homology classes represented by elliptic operators. The precise definition is in Sections 3.3-3.4 via inclusion-exclusion and Kawasaki’s index theorem [11].
We prove the following theorems.
Theorem 1.
(Orbifold version) If $\sigma$ is transverse to the zero section, so that $\mathcal{M}=\sigma^{-1}(0)$ is a stably almost complex sub-orbifold, such extension construction allows us to define the K-theoretic fundamental class $[\mathcal{M}]\in\oplus_{(g)}K_{0}(\Sigma^{N}|\mathcal{M}^{(g)}|^{+})=K_{0}(\mathcal{M})$. Then
$i_{*}[\mathcal{M}]=\mathcal{O}^{vir}_{\mathcal{Y,E},\sigma}.$
Theorem 2.
(Orbifold version) Given a virtual bundle $\mathcal{V}\in K^{0}(\mathcal{Y})$, the pairing of $\mathcal{O}^{vir}_{\mathcal{Y,E},\sigma}$ with $\mathcal{V}$ equals the invariants defined by $\mathcal{V}$ in [2].
Acknowledgments
The author thanks Professor Alexander Givental for drawing his attention to quantum K-theory for Symplectic manifolds, suggesting related problems, and providing patience and guidance throughout the process.
2. Orbifold K-theory
Let $\mathcal{Y}=Y/G$ be a compact orbifold (with boundary), written as a compact manifold $Y$ quotient by a compact Lie group $G$ ([13] Corollary 1.3). We use cursive letters for objects on orbifolds and printed-style letters for $G$-equivariant and $G$-invariant objects on manifolds. We use $\varepsilon$ for the ‘grading’ operator on various $\mathbb{Z}_{2}$-graded objects, i.e., the operator that acts as $(-1)^{a}$ on the subspace with grading $a$. We use $\mathbb{C}(\mathcal{Y})$ for the $\mathbb{C}^{\ast}$-algebra of complex-valued continuous functions (under the $\sup$-norm) on $\mathcal{Y}$.
2.1. (Compactly supported) K-group
We recall the relative K-group $K^{0}(\mathcal{Y},\partial\mathcal{Y})$ of the orbifold $\mathcal{Y}=Y/G$.
For an equivariant bundle map $\alpha:E_{+}\to E_{-}$ between complex $G$-bundles over $Y$, we define the support of $\alpha$ as the (closed) set of points in $Y$ over which $\alpha$ is not an isomorphism of representations of the isotropy groups. The set $\mathcal{C}_{G}(Y,\partial{Y})$ of all bundle map $\alpha$’s supported off $\partial Y$ is a semi-group under direct sum (defined as $\alpha^{0}\oplus\alpha^{1}:E^{0}_{+}\oplus E^{1}_{+}\to E^{0}_{-}\oplus E^{1}_{-}$ for $\alpha^{i}:E^{i}_{+}\to E^{i}_{-}$). The zero element is the identity map between 0-dimensional trivial bundles over $Y$.
An element $\alpha:E_{+}\to E_{-}$ in $\mathcal{C}_{G}(Y,\partial{Y})$ is degenerated if it has empty support. Two elements $\alpha^{0}:E^{0}_{+}\to E^{0}_{-}$ and $\alpha^{1}:E^{1}_{+}\to E^{1}_{-}$ in $\mathcal{C}_{G}(Y)$ are homotopic if there is an element $\tilde{\alpha}:\tilde{E}_{+}\to\tilde{E}_{-}$ in $C_{G}(Y\times[0,1],(\partial Y)\times[0,1])$ that restricts to $\alpha^{0}$ and $\alpha^{1}$ respectively at $Y\times\{0\}$ and $Y\times\{1\}$. Remark that the pair $(Y\times[0,1],(\partial Y)\times[0,1])$ equivariantly retracts to $(Y,\partial Y)$, thus we can take $\tilde{E}_{\pm}$ be the pull-back of a certain bundle over $(Y,\partial Y)$, i.e., independent of $t\in[0,1]$.
We define an equivalence relation on $\mathcal{C}_{G}(Y,\partial{Y})$ generated by identifying degenerate elements with 00 and identifying homotopic elements. We define the $G$-equivariant $K$-group $K_{G}^{0}(Y,\partial{Y})$ of $Y$ as the set of equivalence classes of this equivalence relation under direct sum.
In literature, sometimes we define the K-group $K_{G}^{0}(Y,\partial Y)$ as follows. We define the absolute $K$-groups $K^{0}_{G}$ of a space as the Grothendieck group associated to the semi-group (under direct sum) of isomorphism classes of finite-dimensional complex $G$-bundles over that space. A map between spaces induces a pull-back of bundles, which is a homomorphism between corresponding absolute $K$-groups. We define the reduced $K$-group $\tilde{K}_{G}^{0}(Y/\partial Y,pt)$ as the kernel of the pull-back of $K^{0}_{G}(Y/\partial Y)\to K^{0}_{G}(pt)$ along the inclusion $pt\to Y/\partial Y$ of the base point, and define $K_{G}^{0}(Y,\partial Y)=\tilde{K}_{G}^{0}(Y/\partial Y,pt)$.
Note that by definition, each element in $\tilde{K}_{G}^{0}(Y/\partial Y,pt)$ is represented by some element with $E_{+}|_{U}=E_{-}|_{U}=U\times_{G}V$, for some representation $V$ of $G$, in a $G$-invariant neighborhood $U$ of $\partial Y$. This definition coincides with our first definition using compactly supported bundle homomorphisms ([14] Proposition 3.1, see also [4] page 490). An isomorphism is explicitly constructed as follows. Given $E_{+}\ominus E_{-}\in\tilde{K}_{G}^{0}(Y/\partial Y,pt)$, let $\alpha:E_{+}\to E_{-}$ be an arbitrary map that is the identity map in $U$. The pair $(E_{\pm},\alpha)$ represents an element in $K^{0}_{G}(Y,\partial Y)$. Conversely, given a bundle map $\alpha:E_{+}\to E_{-}$, we pick a bundle $E$ over $Y$, and a $G$-invariant neighborhood $U$ of $\partial Y$, such that $(E_{+}\oplus E)|_{U}$ takes form $U\times_{G}V$. Then $\alpha\oplus id:E_{+}\oplus E\to E_{-}\oplus E$ represents an element in $\tilde{K}_{G}^{0}(Y/\partial Y,pt)$.
For notation simplicity, we define $K^{0}_{G}(Y,\partial{Y})$ as pairs $(E,\alpha)$ of a $\mathbb{Z}_{2}$-graded Hermitian $G$-bundle $E$ and a self-adjoint degree $1$ compactly supported bundle automorphism $\alpha:E\to E$, modulo degenerate elements and homotopy. Indeed, $(E_{+},E_{-},\alpha_{+})$ in the first definition corresponds to $E=E_{+}\oplus E_{-}$ and the map
$\alpha=\left[\begin{matrix}0&\alpha_{+}^{\ast}\\ \alpha_{+}&0\end{matrix}\right]$
Here we identify $(E_{\pm})^{\ast}$ with $E_{\pm}$ by picking a Hermitian metric on $E_{\pm}$. Hermitian metrics on $E_{\pm}$ define homotopic representatives of the $K$-class, thus the same elements in the $K$-group.
The orbifold K group is defined as $K^{0}(Y/G,\partial{Y}/G)=K_{G}^{0}(Y,\partial{Y})$.
2.2. Twisted sectors, inertia orbifold, and fake K-group
Given a quotient orbifold $\mathcal{Y}=Y/G$, we define the inertia orbifold
$I\mathcal{Y}=\left(\coprod_{g\in G}Y^{g}\times\{g\}\right)/G.$
Here $h\in G$ acts on $(x,g)$ as $h\circ(x,g)=(hx,hgh^{-1})$, and $Y^{g}$ is the $g$-fixed locus in $Y$. We define $\mathcal{Y}^{(g)}$ (or simply $\mathcal{Y}^{g}$, if no confusion) as the component in $I\mathcal{Y}$ labeled by the conjugacy class $(g)$. We call $\mathcal{Y}^{(g)}$ sectors and call those with $g\neq 1$ twisted sectors.
More generally, given $\mathbf{g}=(g_{1},\cdots,g_{k})\in G^{k}$, we use $Y^{\mathbf{g}}$ for the locus fixed by $g_{1},\cdots,g_{k}$. We define
$I^{k}\mathcal{Y}=\left(\coprod_{\mathbf{g}}\mathcal{Y}^{\mathbf{g}}\times\{\mathbf{g}\}\right)/G,$
where $h\in G$ acts on $(x,\mathbf{g})$ by $h\circ(x,\mathbf{g})=(hx,h\mathbf{g}h^{-1})$. Define $\mathcal{Y}^{(\mathbf{g})}$ (or $\mathcal{Y}^{\mathbf{g}}$ if no confusion) as the sector in $I^{k}\mathcal{Y}$ corresponding to the conjugacy class $(\mathbf{g})$. We call $\mathcal{Y}^{(\mathbf{g})}$ a $k$-multi-sector.
By ‘fake’ theories of orbifolds, we mean theories in which we only remember the topological space structure on $\mathcal{Y}$ (and forget isotropy groups at points in $\mathcal{Y}$). As an example, define the fake orbifold K-group $K_{\mathbb{Z}}^{0,fake}(\mathcal{Y},\partial\mathcal{Y})$ as classes represented by $(\alpha=\alpha^{\ast}:\mathcal{E}\to\mathcal{E})\in K^{0}(\mathcal{Y},\partial\mathcal{Y})$, with isotropy groups at points in $\mathcal{Y}$ acting trivially on fibers in $\mathcal{E}$. Define
$K^{0,fake}(\mathcal{Y},\partial\mathcal{Y})=K_{\mathbb{Z}}^{0,fake}(\mathcal{Y},\partial\mathcal{Y})\otimes\mathbb{C}.$
Note that the super-trace of a $global$ isotropy $(g)\subseteq G$ defines a map $K^{0}(\mathcal{Y},\partial\mathcal{Y})\otimes\mathbb{C}\to K^{0,fake}(\mathcal{Y},\partial\mathcal{Y})$. Note also that by the eigen-bundle decomposition, there is an isomorphism
$K^{0}(\mathcal{Y},\partial\mathcal{Y})\otimes\mathbb{C}\to K^{0,fake}(I\mathcal{Y},\partial I\mathcal{Y}),$
by sending $\alpha$ to the element with $str_{(g)}i_{(g)}^{\ast}\alpha$ in the twisted sector labeled by $(g)$, where $i_{(g)}:\mathcal{Y}^{(g)}\to\mathcal{Y}$ is the inclusion.
2.3. K-homology
We consider Kasparov $\mathbb{C}(\mathcal{Y})$-modules $(H,\phi,T)$, where by definition $H$ is a $\mathbb{Z}_{2}$-graded complex Hilbert space, and $T$ a degree $1$ self-adjoint automorphism of $H$. The map $\phi:\mathbb{C}(\mathcal{Y})\to L(H)$ is a *-homomorphism from $\mathbb{C}(\mathcal{Y})$ to the space $L(H)$ of bounded linear maps on $H$, such that each $\phi(f)$ is of degree 00 and almost commutes with $T$ in the sense that $[T,\phi(f)]$ is compact in $L(H)$ [9]. The set of Kasparov $\mathbb{C}(\mathcal{Y})$-modules is an abelian semi-group under direct sum: $(H,\phi,T)\oplus(H^{\prime},\phi^{\prime},T^{\prime})=(H\oplus H^{\prime},\phi\oplus\phi^{\prime},T\oplus T^{\prime})$, with the zero element $H=0,\phi(f)=T=0$.
We say that $(H,\phi,T)$ is unitarily equivalent to $(H^{\prime},\phi^{\prime},T^{\prime})$, if there is a degree 00 unitary map $U:H\to H^{\prime}$ that takes $(\phi,T)$ to $(\phi^{\prime},T^{\prime})$. We say that $(H,\phi,T_{0})$ is homotopic to $(H,\phi,T_{1})$ if there is a continuous change of operators $T_{t}:t\in[0,1]$ that connects $T_{0},T_{1}$, such that $(H,\phi,T_{t})$ are Kasparov modules for all $t\in[0,1]$. We say that $(H,\phi,T)$ is degenerate if $T$ is invertible and commutes with $\phi(f)$ for all $f\in\mathbb{C}(\mathcal{Y})$.
We identify Kasparov modules that differ by a degenerated Kasparov $\mathbb{C}(\mathcal{Y}s)$-module and identify unitarily equivalent and homotopic Kasparov modules to obtain the fake $K$-homology (semi-)group $K^{fake}_{0}(\mathcal{Y})$. Indeed, $K^{fake}_{0}(\mathcal{Y})$ is a group [9]. The inverse of $(H,\phi,T)$ is $(H^{\prime},\phi^{\prime},-T)$, where $H^{\prime}$ is $H$ with the grading shifted by $1$, and $\phi^{\prime}=\phi\varepsilon$ with $\varepsilon$ acting as $(-1)^{a}$ on parts with grading $a$ (Proposition 1 in [9]).
We define $K_{0}(\mathcal{Y})=K^{fake}_{0}(I\mathcal{Y})$. Note that a map $f:\mathcal{Y}\to\mathcal{Y}^{\prime}$ induces
$\begin{array}[]{lll}f_{\ast}:&K^{fake}_{0}(\mathcal{Y})&\to K^{fake}_{0}(\mathcal{Y}^{\prime}).\\ &(H,\phi,T)&\to(H,\phi\circ f^{\ast},T)\end{array}$
2.4. Embedding relatively compact sets into compact orbifolds
In this subsection, we prove a technical proposition.
Proposition 2.1.
Any compact subset $\mathcal{U}\subseteq\mathcal{Y}\setminus\partial\mathcal{Y}$ in a Riemannian orbifold $\mathcal{Y}$ has a neighborhood that open embeds into a closed orbifold $\mathcal{W}$.
We first prove the following lemma.
Lemma 2.1.
The subset $\mathcal{U}$ is contained in the interior of a compact sub-orbifold $\mathcal{X}\subseteq\mathcal{Y}\setminus\partial\mathcal{Y}$ whose boundary $\partial\mathcal{X}$ is (an orbifold) transverse to twisted strata.
Proof.
Assume without loss of generality that $\mathcal{U}$ is connected. By compactness of $\mathcal{U}$, there are finitely many geodesic balls $B(x_{i},r_{i})$ that cover $\mathcal{U}$, such that the exponential map at each $x_{i}$ is a diffeomorphism between $B(x_{i},2r_{i})$ and their pre-image. We start with an arbitrary compact co-dimension $1$ orbifold without boundary that meets some $B(x_{i},r_{i})$ and modify this orbifold on all $B(x_{i},r_{i})$’s to obtain a co-dimension $1$ orbifold that bounds a compact region $\mathcal{X}$ which contains all $B(x_{i},r_{i})$.
In each step of modification, we pick a geodesic ball $B(x_{i},r_{i})$ that meets the co-dimension $1$ orbifold and ‘push’ the sub-orbifold (black) out (red) of this ball within $B(x_{i},2r_{i})$, while fixing points near the boundary of $B(x_{i},2r_{i})$ (by pushing or pulling along the radius of corresponding geodesics emerging from $x_{i}$’s). Applying this construction until the resulting co-dimension $1$ orbifold $\partial\mathcal{X}$ does not meet any $B(x_{i},r_{i})$. In this case, $\partial\mathcal{X}$ bounds a compact region $\mathcal{X}$ that contains all $B(x_{i},r_{i})$, thus contains $\mathcal{U}$.
Moreover, since $\mathcal{U}$ is relatively compact, there are finitely many twisted strata. Under the assumption that the space $\mathcal{Y}$ is connected, each stratum is either the entire $\mathcal{Y}$ or a sub-orbifold of strictly lower dimension. We can make $\partial\mathcal{X}$ transverse to these strata by a perturbation. ∎
Proof.
(Proof of Proposition 2.1)
$+$$\mathcal{X}$$\mathcal{X}^{\prime}$$\mathcal{X}$$\mathcal{X}^{\prime}$
There is a collar neighborhood $\varphi:\partial\mathcal{X}\times(0,1)\to\mathcal{X}$ of $\partial\mathcal{X}$ in $\mathcal{X}$. We take two copies $\mathcal{X}=\mathcal{X}^{\prime}$, with collar neighborhoods $\varphi,\varphi^{\prime}$, and define $\mathcal{W}$ as the ‘doubling’
$(\mathcal{X}\coprod\mathcal{X}^{\prime})/(\varphi(x,t)\sim\varphi^{\prime}(x,1-t)).$
Then $\mathcal{W}$ is closed, and there is an embedding $\mathcal{U}\hookrightarrow\mathcal{X}\hookrightarrow\mathcal{W}$. ∎
Corollary 2.1.
Any compact subset $U\subseteq Y\setminus\partial Y$ in a Riemannian $G$-manifold $Y$ has a $G$-invariant neighborhood that equivariantly open embeds into a closed, almost free $G$-manifold $W$.
2.5. The ‘choose-an-operator’ map
For a compact $G$-manifold $Y$ (with boundary), define $T^{\ast}_{G}Y$ as the set of co-vectors on $Y$ annihilating all vectors tangent to $G$-orbits. Add a $G$-invariant Hermitian structure on $Y$ and let $B^{\ast}_{G}Y$ be the unit ball in $T^{\ast}_{G}Y$. In this subsection, we construct a ‘choose-an-operator’ map $\mu=\mu_{G,Y}:K_{G}^{0,inv}(B_{G}^{\ast}Y,\partial B_{G}^{\ast}Y)\to K^{fake}_{0}(Y/G)$, where $K_{G}^{0,inv}$ stands for elements in $K_{G}^{0}$ represented by virtual bundles that correspond to the trivial representation of $G$. The orbifold ‘choose-an-operator’ map $\mu_{Y/G}$ is $\mu_{G,Y}$ sector-wise. The ‘choose-an-operator’ map when $G$ is the trivial group is constructed in, e.g., [5].
We first define certain ‘good’ representatives of elements in $K_{G}^{0,inv}(B_{G}^{\ast}Y,\partial B_{G}^{\ast}Y)$, which we call symbolic representatives, that serve as symbols of certain pseudo-differential operators representing K-homology classes.
If any two automorphisms of the same bundle agree outside a compact set, they represent the same element in the $K$ group. So to describe an element in $K_{G}^{0,inv}(B_{G}^{\ast}Y,\partial B_{G}^{\ast}Y)$, we only need to define the bundle involved and the automorphism near the boundary. Let $\pi_{Y}:B_{G}^{\ast}Y\to Y$, and let $p:\pi_{Y}^{\ast}E\to\pi_{Y}^{\ast}E$ be a bundle isomorphism, defined outside a compact subset which we refer to as $\operatorname{supp}p$, in $Y\setminus\partial Y\subseteq B_{G}^{\ast}Y$.
$Y$$B_{G}^{\ast}Y$$supp\ p$Symbolic representative $p$:Constant along each vertical line segment in $B_{G}^{\ast}Y\setminus supp\ p$
A symbolic representative $p$ supported on $\operatorname{supp}p\subseteq B_{G}^{\ast}(Y)$ is an automorphism of $\pi_{Y}^{\ast}E\setminus E|_{\operatorname{supp}p}$, with
  • $E$ a $G$-bundle over $Y$ that corresponds to the trivial representation and is trivial outside $\operatorname{supp}p$
  • $p(y,\xi)=p(y,\nu\xi)$ for all $(y,\xi)\in B_{G}^{\ast}Y\setminus Y$ and $\nu\in\mathbb{R}_{+}$.
Symbolic representatives modulo certain relations (homotopy and isomorphism) comprise a group that is isomorphic to $K_{G}^{0,inv}(B_{G}^{\ast}Y,\partial B_{G}^{\ast}Y)$ (cf. [4] page 491-493, also [3] page 15). Moreover, we can take this symbolic representative to be self-adjoint.
By Corollary 2.1, we can equivariantly embed a $G$-invariant neighborhood of $\operatorname{supp}p$ into a closed manifold $W$ with almost free $G$-action. Let $E_{\pm}$ be the puls and minus part of the $\mathbb{Z}_{2}$ grading on $E$, and $p_{+}=p|_{E_{+}}$. There are sections of $E_{\pm}|_{Y\setminus\operatorname{supp}p}$ that serve as a basis on each fiber, under which the matrix of $p_{+}$ is the identity matrix. Thus we can extend $p,E$ trivially to $p_{W},E_{W}$ on $W$. Let $P_{W}:C^{\infty}(W,E_{W})\to C^{\infty}(W,E_{W})$ be the quantization of the (self-adjoint) symbolic representative $p_{W}$. Then $P_{W}$ is a self-adjoint pseudo-differential operator of order 00, and is of degree $1$ under the $\mathbb{Z}_{2}$-grading. If $s\in C^{\infty}(W,E_{W})$ is supported in $U\supseteq\operatorname{supp}p$, then $P_{W}s$ is also supported in $U$. If $s\in C^{\infty}(W,E_{W})$ is supported in $W\setminus\operatorname{supp}p$, then $P_{W}s=s$. Let $H=L^{2}(U,E|_{U})^{G}$ be the space of $G$-invariant $L^{2}$-sections of $E$ restricted to $U$, which is identified with a subset of $L^{2}(W,E_{W})$ by extending each section by 00 outside $U$. Since $P_{W}:C^{\infty}(W,E_{W})\to C^{\infty}(W,E_{W})$ is elliptic, it uniquely extends to $L^{2}$-sections in a continuous way. Let $T$ be the restriction to subspace $H$ of this continuous ($G$-equivariant) extension. Let ${\phi}$ be the $\mathbb{C}(Y/G)$-module structures on ${H}$, given by multiplying (the restriction to $U$ of) a continuous $G$-invariant function. We define
$\mu_{G,Y}(\pi_{Y}^{\ast}E,p)=(H,\phi,T).$
Lemma 2.2.
The element $({H},{\phi},{T})$ represents an element in $K^{fake}_{0}(\mathcal{Y})$.
Proof.
Following [3], we show that $T$ is Fredholm by showing that the $G$-invariant parts $(\ker P_{W})^{G}=(\ker P_{W}^{\ast})^{G}$ are finite-dimensional. Pick a bi-invariant Riemannian metric on $G$, and let $v_{1},\cdots,v_{k}$ be first-order differential operators acting on sections of $E$ corresponding to some orthonormal basis of the Lie algebra of $G$. Consider the bundle map $\pi_{W}^{\ast}E\to\pi_{W}^{\ast}E$ given by
$\delta_{G}:(x,\xi,e)\to\left(x,\xi,\frac{\sum_{i}\left<v_{i},\xi\right>^{2}}{1+\sum_{i}\left<v_{i},\xi\right>^{2}}e\right).$
The quantization of $\delta_{G}$ gives an order 00 pseudo-differential operator whose kernel contains the space of $G$-invariant functions as a closed subspace. Now, the operator $A=(P_{W},\Delta_{G})$ has an injective symbol, so the symbol of $A^{\ast}A$ is invertible. Thus the operator $A^{\ast}A=P_{W}^{\ast}P_{W}+\Delta^{\ast}_{G}\Delta_{G}$ is elliptic with finite-dimensional kernel. Now $(\ker P_{W})^{G}=(\ker P^{\ast}_{W})^{G}$ is contained in the kernel of $A^{\ast}A$, so it is also finite-dimensional. Thus $T$ is Fredholm.
Moreover, for an arbitrary $f$, the order 00 symbol of the commutator $[P_{W}|_{U},\phi(f)]$ is 00, so $[T,\phi(f)]$ is compact. Thus $\mu_{G,Y}(\pi_{W}^{\ast}E_{W},p_{W})\in K^{fake}_{0}(Y)$. ∎
Proposition 2.2.
The operator $\mu_{G,Y}$ is well-defined as a map $K_{G}^{0,inv}(B_{G}^{\ast}Y,\partial B_{G}^{\ast}Y)\to K_{0}^{fake}(Y)$.
Proof.
We temporarily denote the ‘choose-an-operator’ map with choice $U,W,p$ by $\mu_{G,Y}^{U,W}(p)$. We need to prove:
  1. (1)
    $\mu_{G,Y}^{U,W}(p)$ is independent of the symbolic representative $p$.
  2. (2)
    $\mu_{G,Y}^{U,W}([p])$ is independent of $U,W$.
For fixed $U,W$, let $P_{W},P_{W}^{\prime}$ be two pseudo-differential operators with symbolic representatives $[p_{W}]=[p_{W}^{\prime}]$. If $p_{W},p_{W}^{\prime}$ are homotopic (in particular, if $p_{W}=p_{W}^{\prime}$ but $P_{W}\neq P_{W}^{\prime}$), i.e., their symbols are connected by a homotopy of degree $1$ automorphisms of the $same$ bundle $\pi^{\ast}_{W}E$, then the Hilbert spaces involved are the same, and the operators are connected by a homotopy of operators given by the quantization of the symbols, thus connected by a homotopy of representatives of $K$-homology. If the $p_{W}$ is degenerated $K_{G}^{0,inv}(B_{G}^{\ast}W,\partial B_{G}^{\ast}W)$, then it is an isomorphism between positive and negative parts in the $\mathbb{Z}_{2}$-grading of the ambient bundle with empty support. Computing under the identification given by this isomorphism, the operator $T$ is identity, and thus $\mu_{G,Y}(p_{W})$ is degenerated.
For fixed $U$, the Hilbert spaces are naturally identified, and different $W$’s give the same operator $T$. For $\operatorname{supp}p\subseteq U^{\prime}\subseteq U$, let $(E,p)\in K_{G}^{0,inv}(B_{G}^{\ast}Y,\partial B_{G}^{\ast}Y)$ be a symbolic representative. The corresponding $K$-homology element defined by $U^{\prime}$ and $U$ differs by the image of the degenerated element $(E,p)|_{U\setminus U^{\prime}}$ under the choose-an-operator map, thus they define the same element in $K_{0}^{fake}(Y)$. Two general compact sets are contained in their union, so they also define the same element in $K_{0}^{fake}(Y)$. ∎
We have the following direct corollary.
Corollary 2.2.
The choose-an-operator map is compatible with open embeddings.
The cotangent bundle of $\mathcal{Y}=Y/G$ is $T^{\ast}\mathcal{Y}=(T^{\ast}_{G}Y)/G$. Denote its unit ball bundle by $B^{\ast}\mathcal{Y}$ (under a Hermitian metric). Observe that the sector $(B^{\ast}\mathcal{Y})^{(g)}=B^{\ast}(\mathcal{Y})^{(g)}$, and that the supertrace $str_{g}$ maps $K^{0,fake}(B^{\ast}\mathcal{Y}^{(g)},\partial B^{\ast}\mathcal{Y}^{(g)})$ to $K_{G}^{0,inv}(B_{G}^{\ast}Y^{(g)},\partial B_{G}^{\ast}Y^{(g)})\otimes\mathbb{C}$. We define the ‘genuine’ choose-an-operator map $\mu$ sector-wise, as the invariant choose-an-operator map composed with supertraces. The orbifold choose-an-operator map is independent of the presentation of the orbifold.
3. The pairing
3.1. The Cap product
In this Subsection, we define the K-theoretic cap product on $\mathcal{Y}$. Given $(H,\phi,T)\in K^{fake}_{0}(\mathcal{Y})$ and $(\mathcal{E},\alpha)\in K^{0,fake}(\mathcal{Y},\partial\mathcal{Y})$, we construct an element $(\tilde{H},\tilde{\phi},\tilde{T})\in K^{fake}_{0}(\mathcal{Y})$ as the cap product $\alpha\cap T$ following [8].
We first describe the Hilbert $\mathbb{C}(\mathcal{Y})$-module structure $(\tilde{H},\tilde{\phi})$. For a complex orbi-bundle $\mathcal{E}$, let $C(\mathcal{E})$ be continuous sections of $\mathcal{E}$. Then $C(\mathcal{E})$ is a finitely generated projective module over $\mathbb{C}(\mathcal{Y})$. So there is an idempotent matrices $A$ of size $n$, with entries in $\mathbb{C}(\mathcal{Y})$, whose range is $C(\mathcal{E})$. It is viewed as a map $A:H^{n}\to H^{n}$ via $\phi:\mathbb{C}(\mathcal{Y})\to L(H)$. Set $\tilde{H}=\operatorname{im}A$ and $\tilde{\phi}$ be the natural $\mathbb{C}(\mathcal{Y})$-module structures on $\tilde{H}$.
To describe the Fredholm operator $\tilde{T}:\tilde{H}\to\tilde{H}$, let $C(\alpha)$ be the $\mathbb{C}(\mathcal{Y})$-module homomorphism $C(\mathcal{E})\to C(\mathcal{E})$ induced by $\alpha:\mathcal{E}\to\mathcal{E}$. Let $T\otimes^{\prime}1,1\otimes C(\alpha)$ be the induced maps on $\operatorname{im}A$, and
$\tilde{T}=T\otimes^{\prime}1-\varepsilon 1\otimes C(\alpha).$
We give the following remarks.
  1. (1)
    We have $\tilde{H}=H\otimes_{\mathbb{C}(\mathcal{Y})}C(\mathcal{E})$ as vector spaces, but different $A$’s define on this vector space a different inner product. Nonetheless, all such inner products define the same topology.
  2. (2)
    While $1\otimes C(\alpha)$ is the actual tensor product, we emphasize that the ‘tensor product’ $T\otimes 1$ does not make sense, since the (pseudo-differential) operator $T$ is not $\mathbb{C}(\mathcal{Y})$-linear. Therefore, we use the notation $\otimes^{\prime}$ and need to introduce the projection $A$ in our construction.
Lemma 3.1.
The element $(\tilde{H},\tilde{\phi},\tilde{T})$ represents an element in $K^{fake}_{0}(\mathcal{Y})$.
Proof.
Observe that $T\varepsilon+\varepsilon T=0,\varepsilon^{2}=1$, and that $C(\alpha)$ commutes with $\tilde{\phi}(f)$, so we have
$\begin{array}[]{ll}[\tilde{T},\tilde{\phi}(f)]&=[T,\phi(f)]\otimes^{\prime}1;\\ \tilde{T}^{2}&=T^{2}\otimes^{\prime}1+1\otimes C(\alpha)^{2}.\end{array}$
Hence $\tilde{T}$ almost commutes with $\tilde{\phi}(f)$ and $\tilde{T}^{\ast}=\tilde{T}$ has finite dimensional kernel. ∎
Lemma 3.2.
$(H,\phi,T)\in K^{fake}_{0}(\mathcal{Y})$ only depends on $T,\phi$ and the topology on $H$ (and is independent of the inner products on $H$).
Proof.
Let $(H,\phi,T)$ be a Kasparov module. Let $H^{\prime}=H$ as $\mathbb{C}(\mathcal{Y})$-modules and topological spaces, each with a certain inner product. Then
$\tilde{T}_{t}=\left[\begin{matrix}T\cos(\frac{\pi}{2}t)&\sin(\frac{\pi}{2}t)I\\ \sin(\frac{\pi}{2}t)I&-T^{\prime}\cos(\frac{\pi}{2}t)\end{matrix}\right]$
defines a homotopy from $(H,\phi,T)\ominus(H^{\prime},\phi,T)$ to a degenerate Kasparov module. So $(H^{\prime},\phi,T)=(H,\phi,T)$ in $K^{fake}_{0}(\mathcal{Y})$. ∎
Proposition 3.1.
The (fake) cap product is a map $K_{0}^{fake}(\mathcal{Y})\otimes K^{0,fake}(\mathcal{Y},\partial\mathcal{Y})\to K_{0}^{fake}(\mathcal{Y})$.
Proof.
Given $(H,\phi,T)\in K_{0}^{fake}(\mathcal{Y})$ and $(\mathcal{E},\alpha)\in K^{0,fake}(\mathcal{Y},\partial\mathcal{Y})$, we show that
  1. (1)
    The cap product $(\tilde{H},\tilde{\phi},\tilde{T})$ is independent of $A$ (and $n$) chosen.
  2. (2)
    The cap product is independent of the representatives $(H,\phi,T)$ and $(\mathcal{E},\alpha)$ chosen.
Let $A,A^{\prime}$ be idempotent matrices of sizes $n,n^{\prime}$ with $\mathbb{C}(\mathcal{Y})$ coefficients, whose images are identified via a $\mathbb{C}(\mathcal{Y})$-module homomorphism $\iota$ that is also an isometry. Then the composition
[…]
is a lift of $\iota$ as a module homomorphism, thus realizing the map $\tilde{H}\to\tilde{H}^{\prime}$ as a matrix transformation, and so this map is continuous. Likewise the map $\tilde{H}^{\prime}\to\tilde{H}$ induced by $\iota^{-1}$ is also continuous, so the topologies on $\tilde{H}$ and $\tilde{H}^{\prime}$ are identical. Moreover, $\tilde{\iota}$ intertwines $\phi(f)$’s and $T$’s up to a compact operator, so different $A$’s give unitarily equivalent cap products.
Different representatives of $K$-homology differ by a sequence of unitary equivalence, homotopy, or direct sum with degenerate objects. Different representatives of the same element in the $K$ group differ by a sequence of isomorphism, homotopy, or direct sum with degenerated objects. The cap product, in turn, is also connected by a sequence of unitary equivalence, homotopy, or direct sum with degenerated objects. So the cap product
$\cap:K^{fake}_{0}(\mathcal{Y})\otimes K^{0,fake}(\mathcal{Y},\partial\mathcal{Y})\to K^{fake}_{0}(\mathcal{Y})$
is well-defined. ∎
Lemma 3.3.
The (fake) cap product is compatible with embeddings.
Proof.
Given an inclusion of compact sets $\mathcal{Y}\subseteq\mathcal{Y}^{\prime}$, we viewed $C(\mathcal{E}|_{\mathcal{Y}})$ as the range of a $\mathbb{C}(\mathcal{Y})$-matrix. Indeed, restricting the coefficients of this matrix (that are functions on $\mathcal{Y}^{\prime}$) to $\mathcal{Y}$, we obtain a matrix whose range gives $C(\mathcal{E}|_{\mathcal{Y}})$. Computing under this specific choice of basis shows that each inclusion of $\mathcal{Y}$ induces a map on cap products. ∎
The genuine cap product
$\cap:K_{0}(\mathcal{Y})\otimes K^{0}(\mathcal{Y},\partial\mathcal{Y})\to K_{0}(\mathcal{Y})$
is defined sector-wise.
3.2. $K^{0}(\mathcal{Y},\partial\mathcal{Y})$-module structures
In this Subsection, we show that the $K^{0}(\mathcal{Y},\partial\mathcal{Y})$ module structures on $K^{0}(B^{\ast}\mathcal{Y},\partial B^{\ast}\mathcal{Y})$ and $K_{0}(\mathcal{Y})$ are intertwined by $\mu_{\mathcal{Y}}$.
Let $\pi:B^{\ast}\mathcal{Y}\to\mathcal{Y}$. Given elements $(\mathcal{E},\alpha)\in K^{0,fake}(\mathcal{Y},\partial\mathcal{Y})$ and $(\mathcal{F},\beta)\in K^{0,fake}(B^{\ast}\mathcal{Y},\partial B^{\ast}\mathcal{Y})$, we define ([4] page 490)
$\alpha\boxtimes\beta=\beta\otimes 1-\varepsilon\otimes\pi^{\ast}\alpha:\mathcal{F}\otimes\pi^{\ast}\mathcal{E}\to\mathcal{F}\otimes\pi^{\ast}\mathcal{E}.$
This element is supported in the (compact) intersection of $\operatorname{supp}\alpha\subseteq\mathcal{Y}\subseteq B^{\ast}\mathcal{Y}$ and $\operatorname{supp}\beta\subseteq B^{\ast}\mathcal{Y}$, making $K^{0,fake}(B^{\ast}\mathcal{Y},\partial B^{\ast}\mathcal{Y})$ a $K^{0,fake}(\mathcal{Y},\partial\mathcal{Y})$ module. On the other hand, $K^{fake}_{0}(\mathcal{Y})$ carries a $K^{0,fake}(\mathcal{Y},\partial\mathcal{Y})$-module structure given by the cap product. We use $\hat{\otimes}$ for the completion of the algebraic tensor product.
Lemma 3.4.
Let $\mathcal{E},\mathcal{F}$ be Hermitian bundles with trivial isotropy group action over a compact orbifold $\mathcal{Y}$. Then $L^{2}(\mathcal{E})\hat{\otimes}_{\mathbb{C}(\mathcal{Y})}C(\mathcal{F})\cong L^{2}(\mathcal{E}\otimes_{\mathbb{C}}\mathcal{F})$ as $\mathbb{C}(\mathcal{Y})$-modules and as topological spaces.
Proof.
If $\mathcal{F}=\underline{n}$, the $n$-dimensional trivial bundle over $\mathcal{Y}$, $C(\underline{n})=(\mathbb{C}(\mathcal{Y}))^{n}$ induces an isomorphism
$L^{2}(\mathcal{E})\hat{\otimes}_{\mathbb{C}(\mathcal{Y})}C(\underline{n})=(L^{2}(\mathcal{E})\hat{\otimes}_{\mathbb{C}(\mathcal{Y})}\mathbb{C}(\mathcal{Y}))^{n}=(L^{2}(\mathcal{E}\otimes\underline{1}))^{n}=L^{2}(\mathcal{E}\otimes\underline{n}),$
as $\mathbb{C}(\mathcal{Y})$-modules and topological spaces.
For general $\mathcal{F}$, there is an $n$ and a complex vector bundle $\mathcal{F}^{\prime}$ such that $\mathcal{F}\oplus\mathcal{F}^{\prime}=\underline{n}$ by the compactness of $\mathcal{Y}$. The natural map $L^{2}(\mathcal{E})\hat{\otimes}_{\mathbb{C}(\mathcal{Y})}C(\mathcal{F})\to L^{2}(\mathcal{E}\otimes\underline{n})$ is injective with image contained in $L^{2}(\mathcal{E}\otimes\mathcal{F})$, since $C(\underline{n})=C(\mathcal{F})\oplus C(\mathcal{F}^{\prime})$. So $L^{2}(\mathcal{E})\hat{\otimes}_{\mathbb{C}(\mathcal{Y})}C(\mathcal{F})$ is viewed as a sub-module of $L^{2}(\mathcal{E}\otimes\mathcal{F})$. Similarly $L^{2}(\mathcal{E})\hat{\otimes}_{\mathbb{C}(\mathcal{Y})}C(\mathcal{F}^{\prime})\subseteq L^{2}(\mathcal{E}\otimes\mathcal{F}^{\prime})$. Now, it follows from
$L^{2}(\mathcal{E})\hat{\otimes}_{\mathbb{C}(\mathcal{Y})}C(\mathcal{F})\oplus L^{2}(\mathcal{E})\hat{\otimes}_{\mathbb{C}(\mathcal{Y})}C(\mathcal{F}^{\prime})=L^{2}(\mathcal{E}\otimes\underline{n})=L^{2}(\mathcal{E}\otimes\mathcal{F})\oplus L^{2}(\mathcal{E}\otimes\mathcal{F}^{\prime})$
that
$L^{2}(\mathcal{E})\hat{\otimes}_{\mathbb{C}(\mathcal{Y})}\mathbb{C}(\mathcal{F})=L^{2}(\mathcal{E}\otimes\mathcal{F}).$
It is also easily seen from the proof that the topologies are preserved. ∎
Proposition 3.2.
The following diagram commutes.
[…]
Proof.
We only need to prove the sector-wise commutativity
[…]
To do this, let $\beta:\mathcal{F}\to\mathcal{F}$ be an element in $K^{0,fake}(\mathcal{Y})$. Let $P:L^{2}(\mathcal{E})\to L^{2}(\mathcal{E})$ be a pseudo-differential operator of order 00 that represents $\alpha:\mathcal{E}\to\mathcal{E}$. Let $\mathcal{U}$ be a relatively compact open sub-manifold, such that $\alpha$ is an isomorphism, and $P$ is multiplication by a function outside $\mathcal{U}$. We have the following results.
$\begin{array}[]{lll}\beta\boxtimes\sigma(P)&=\beta\otimes 1-\varepsilon\otimes\sigma(P):&\pi^{\ast}\mathcal{E}\otimes\pi^{\ast}\mathcal{F}\to\pi^{\ast}\mathcal{E}\otimes\pi^{\ast}\mathcal{F}\\ P\cap\beta&=P\otimes 1-\varepsilon\otimes C(\beta):&L^{2}(\mathcal{E})\hat{\otimes}_{\mathbb{C}(\mathcal{Y})}\mathbb{C}(\mathcal{F})\to L^{2}(\mathcal{E})\hat{\otimes}_{\mathbb{C}(\mathcal{Y})}\mathbb{C}(\mathcal{F})\end{array}$
The operator
$[P]\cap\beta=[P]\otimes 1-\varepsilon\otimes C(\beta):L^{2}(\mathcal{E}\otimes\mathcal{F})\to L^{2}(\mathcal{E}\otimes\mathcal{F})$
represents $\beta\boxtimes\sigma(P)$ because $\sigma(P\otimes 1)=\sigma(P)\otimes 1$. It also represents $[P]\cap\beta$ by the previous lemma. Therefore $\mu([\sigma(P)]\boxtimes[\beta])=[P]\cap[\beta]$, and so the diagram commute. ∎
3.3. K-homological maps $i^{!}$ and $i_{\ast}$
3.3.1. Statement of results
Let $i:\mathcal{Z}\to\mathcal{Y}$ be a closed suborbifold of $\mathcal{Y}$, whose boundary is contained in that of $\mathcal{Y}$. Let $\pi:\mathcal{N}\to\mathcal{Z}$ be a closed tubular neighborhood with $\pi^{-1}(\partial\mathcal{Z})\subseteq\partial\mathcal{Y}$. Define various projections by the following diagram.
${B^{\ast}\mathcal{N}}$${B^{\ast}\mathcal{Z}}$${\mathcal{N}}$${\mathcal{Z}}$$\scriptstyle{\tilde{\pi}}$$\scriptstyle{\pi_{\mathcal{N}}}$$\scriptstyle{\pi_{\mathcal{Z}}}$$\scriptstyle{\pi}$
Recall that $i_{\ast}^{fake}(H,\phi,T)=(H,\phi\circ i^{\ast},T)$.
Given an inclusion $j:\mathcal{N}\hookrightarrow\mathcal{Y}$, let $\mathcal{I}$ be the ideal of functions in $\mathbb{C}(\mathcal{Y})$ that vanish on $\mathcal{N}$. Given an element $(H,\phi,T)\in K_{0}^{fake}(\mathcal{Y})$, let $H_{\mathcal{N}}=H/\mathcal{I}H$, $\phi_{\mathcal{N}}(f)$ the action of an arbitrary extension of $f$ on the quotient space $H_{\mathcal{N}}$ via $\phi$, and $T_{\mathcal{N}}:H_{\mathcal{N}}\to H_{\mathcal{N}}$ the quotient map of $T$. We define
$\begin{array}[]{llll}j^{!}:&K_{0}^{fake}(\mathcal{Y})&\to&K_{0}^{fake}(\mathcal{N})\\ &(H,\phi,T)&\to&(H_{\mathcal{N}},\phi_{\mathcal{N}},T_{\mathcal{N}})\end{array}$
and $i^{fake,!}$ as the composition $\pi_{\ast}j^{!}$.
We also recall the wrong-way $K$-cohomological push-forward
$i_{!}^{fake}:K^{0,fake}(B^{\ast}\mathcal{Z},\partial B^{\ast}\mathcal{Z})\to K^{0,fake}(B^{\ast}\mathcal{N},\partial B^{\ast}\mathcal{N})$
defined in [4]. Let $\mathcal{N}_{\mathbb{C}}=B(\pi_{\mathcal{Z}}^{\ast}\mathcal{N}\otimes\mathbb{C})$, which is identified with $B^{\ast}\mathcal{N}\to B^{\ast}\mathcal{Z}$ by identifying the real part with fibers of $\mathcal{N}\to\mathcal{Z}$ and imaginary part with corresponding cotangent vectors. We define the K-theoretical push-forward in the special case of $\mathcal{Y}=\mathcal{N}$ as
$i_{!}^{fake}:\alpha\to(\pi^{\ast}\alpha)\otimes(str_{1}\wedge^{\ast}\mathcal{N}_{\mathbb{C}}^{\ast})$
The push-forward $i_{!}^{fake}$ in the general case is the composition of $i_{!}^{fake}$ for tubular neighborhood and the map $K^{0,fake}(B^{\ast}\mathcal{N},\partial B^{\ast}\mathcal{N})\to K^{0,fake}(B^{\ast}\mathcal{Y},\partial B^{\ast}\mathcal{Y})$ given by extending a virtual bundle trivially outside $\mathcal{N}$. $i_{!}^{fake}$ is independent of the tubular neighborhood chosen [3].
We define the genuine $i_{!},i_{\ast},i^{!}$ sector-wise, and will show in this subsection that
Proposition 3.3.
  1. (1)
    The homological $i^{!}$ is well defined, and the following diagram commutes.
    ${K^{0}(B^{\ast}\mathcal{Z},\partial B^{\ast}\mathcal{Z})}$${K^{0}(B^{\ast}\mathcal{Y},\partial B^{\ast}\mathcal{Y})}$${K_{0}(\mathcal{Z})}$${K_{0}(\mathcal{Y})}$$\scriptstyle{i_{!}}$$\scriptstyle{\mu_{\mathcal{Z}}}$$\scriptstyle{\mu_{\mathcal{Y}}}$$\scriptstyle{i^{!}}$
  2. (2)
    The following diagram commutes.
    ${K^{0}(B^{\ast}\mathcal{Z},\partial B^{\ast}\mathcal{Z})}$${K^{0}(B^{\ast}\mathcal{Y},\partial B^{\ast}\mathcal{Y})}$${K_{0}(\mathcal{Z})}$${K_{0}(\mathcal{Y})}$$\scriptstyle{i_{!}}$$\scriptstyle{\mu_{\mathcal{Z}}}$$\scriptstyle{\mu_{\mathcal{Y}}}$$\scriptstyle{i_{\ast}}$
3.3.2. A technical lemma
Our goal in this Subsection is to prove
Lemma 3.5.
The following diagram commutes.
${K^{0,fake}(B^{\ast}\mathcal{Z},\partial B^{\ast}\mathcal{Z})}$${K^{0,fake}(B^{\ast}\mathcal{N},\partial B^{\ast}\mathcal{N})}$${K^{fake}_{0}(\mathcal{Z})}$${K^{fake}_{0}(\mathcal{N})}$$\scriptstyle{i_{!}^{fake}}$$\scriptstyle{\mu_{\mathcal{Z}}}$$\scriptstyle{\mu_{\mathcal{N}}}$$\scriptstyle{\pi_{\ast}}$
We first fix some notations. Fix a symbolic representative $(\pi_{\mathcal{Z}}^{\ast}\mathcal{E},\alpha)\in K^{0,fake}(B^{\ast}\mathcal{Z},\partial B^{\ast}\mathcal{Z})$. We define $\beta$ and $\tilde{D}$, and a specific $\tilde{T}$ (and $T$) that fit into the following diagram. Note that the ‘product’ $\otimes^{\prime}$ is as defined in Section 3.1 and is not the genuine tensor product, as $\tilde{T}$ and $\tilde{D}$ are not $\mathbb{C}(\mathcal{N})$-linear. Nonetheless, $\otimes^{\prime}$ is the genuine tensor product $\otimes$ after pushing-forward to $K_{0}^{fake}(\mathcal{Z})$ under $\pi_{\ast}$.
${(\pi_{\mathcal{Z}}^{\ast}\mathcal{E},\alpha)}$${(\tilde{\pi}^{\ast}\pi^{\ast}_{\mathcal{Z}}\mathcal{E}\otimes str_{1}\wedge^{\ast}\mathcal{N}_{\mathbb{C}}^{\ast},\alpha\otimes 1-\varepsilon 1\otimes\beta)}$${(L^{2}(\mathcal{Z},\mathcal{E}),T)}$${(L^{2}(\mathcal{N},\pi^{\ast}(\mathcal{E}\otimes str_{1}(\wedge^{\ast}\mathcal{N}^{\ast}\otimes\mathbb{C}))),\tilde{T}\otimes^{\prime}1-\varepsilon 1\otimes^{\prime}\tilde{D})}$$\scriptstyle{i_{!}^{fake}}$$\scriptstyle{\mu_{\mathcal{Z}}}$$\scriptstyle{\mu_{\mathcal{N}}}$
Let $\beta$ be a symbolic representative of $\lambda_{\xi}:\wedge^{\ast}\mathcal{N}_{\mathbb{C}}\to\wedge^{\ast}\mathcal{N}_{\mathbb{C}}$ (as defined in the introduction) in $K^{0,fake}(\mathcal{N})$ and $\tilde{D}$ its quantization. Then $\tilde{D}$ is a self-adjoint 0-th order pseudo-differential operator over $\mathcal{N}$ that acts in the fiber direction of $\mathcal{N}\to\mathcal{Z}$.
We write $\mathcal{Z}=Z/G,\mathcal{N}=N/G$ and pick an equivariant, order 00 pseudo-differential operator $P$ on $Z$ whose symbol represents $\alpha$. Recall that in Section 2.5, we defined $T$ as (the average of) $P$ restricted to $G$-invariant $L^{2}$-sections. To define $\tilde{T}$, we pick $\mathcal{N}^{\prime}$ such that $\mathcal{N}\oplus\mathcal{N}^{\prime}=Z\times_{G}V$ for $V$ (the unit ball in) some $G$-representation. We construct $\tilde{P}$ as the order 00 pseudo-differential operator on $Z\times V$ given by $P$ on each slice $Z\times\{v\}$, and $\tilde{P}^{G}$ as the average of $\tilde{P}$ over $G$. We define $\tilde{T}$ by its action on a section of $\pi^{\ast}\mathcal{E}$ over $\mathcal{N}$ equals the restriction to $\mathcal{N}$ of $\tilde{P}_{G}$ acting on the pull-back of that section along $\mathcal{N}\oplus\mathcal{N}^{\prime}\to\mathcal{N}$.
Lemma 3.6.
For an element $(\mathcal{H},\phi,\mathcal{T})\in K^{fake}_{0}(\mathcal{Z})$, let $\Delta\in L(\mathcal{H})$ be a degree $1$ self adjoint operator that commutes with $\phi$ and $\mathcal{T}$. Then $(\mathcal{H},\phi,\mathcal{T}-\varepsilon\Delta)=(\ker\Delta,\phi|_{\ker\Delta},\mathcal{T})$.
Proof.
The operators $\mathcal{T}$ and $\Delta$ commute, so $\ker\Delta$ and $\operatorname{im}\Delta$ are both invariant under $\mathcal{T}-\varepsilon\Delta$ and $\phi(f)$. Thus $(\ker\Delta,\phi|_{\ker\Delta},\mathcal{T})$ and $(\operatorname{im}\Delta,\phi|_{\operatorname{im}\Delta},\mathcal{T}-\varepsilon\Delta)$ both represent elements in $K_{0}^{fake}(\mathcal{Z})$, and
$(\ker\Delta,\phi|_{\ker\Delta},\mathcal{T})\oplus(\operatorname{im}\Delta,\phi|_{\operatorname{im}\Delta},\mathcal{T}-\varepsilon\Delta)=(\mathcal{H},\phi,\mathcal{T}-\varepsilon\Delta).$
Now $(\operatorname{im}\Delta,\phi|_{\operatorname{im}\Delta},\mathcal{T}-\varepsilon\Delta)$ is homotopic to the degenerated element $(\operatorname{im}\Delta,\phi|_{\operatorname{im}\Delta},-\varepsilon\Delta)$ by the line segment connecting them, so
$(\ker\Delta,\phi|_{\ker\Delta},\mathcal{T})=(\mathcal{H},\phi,\mathcal{T}-\varepsilon\Delta).$
Proof.
(Proof of Lemma 3.5)
Observe that the image under $\pi_{\ast}$ of $1\otimes\tilde{D}$ commutes with the images of $\tilde{T}\otimes 1$ and $\phi$, so by the previous lemma we know that $\pi_{\ast}\mu_{\mathcal{N}}i_{!}^{fake}(\mathcal{E},\alpha)=(\ker 1\otimes\tilde{D},\tilde{T}\otimes 1|_{\ker 1\otimes\tilde{D}})$.
Let $s_{i}$’s be a $\mathbb{C}(\mathcal{Z})$-basis of $C(\mathcal{N})$, and form $\beta_{i},\tilde{D}_{i}$ from $s_{i}$’s as we form $\beta,\tilde{D}$. Then $1\otimes\tilde{D}_{i}$ commutes with $\tilde{T}\otimes 1$ and $\phi$. If $(V,\tilde{T}\otimes 1)\in K_{0}^{fake}(\mathcal{Z})$, then the self-adjoint operator $\tilde{T}\otimes 1$ almost commutes with the $\mathbb{C}(\mathcal{Z})$-action, and so represents the same element as $(V,\tilde{T}\otimes 1-\varepsilon\tilde{D}_{i})$. Thus, again, by the previous lemma, we can restrict $V$ to $\ker\tilde{D_{i}}$. So $\pi_{\ast}\mu_{\mathcal{N}}i_{!}^{fake}(\mathcal{E},\alpha)=(\cap_{i}\ker(1\otimes\tilde{D}_{i}),\tilde{T}\otimes 1)$.
On each fiber of $\mathcal{N}\to\mathcal{Z}$, the equations given by $\tilde{D}_{i}=0$ precisely determine a section over $\mathcal{N}$ from its initial values in $\mathcal{Z}$ ([4]). Thus the common kernels of $\tilde{D}_{i}$ is isomorphic to $(L^{2}(\mathcal{Z},\mathcal{E}),\tilde{T}\otimes 1)=(L^{2}(\mathcal{E}),T\otimes 1)$, and the proposition follows. ∎
3.3.3. Proof of Proposition 3.3
It suffices to prove the statement sector-wise, and that the statement on twisted sectors is identical to that on the non-twisted sector (only with $\mathcal{Z},\mathcal{N},\mathcal{Y}$’s replaced by the corresponding twisted sectors in these spaces).
Corollary 3.1.
The following diagram commutes.
${K^{0,fake}(B^{\ast}\mathcal{Z},\partial B^{\ast}\mathcal{Z})}$${K^{0,fake}(B^{\ast}\mathcal{N},\partial B^{\ast}\mathcal{N})}$${K^{fake}_{0}(\mathcal{Z})}$${K^{fake}_{0}(\mathcal{N})}$$\scriptstyle{i_{!}^{fake}}$$\scriptstyle{\mu_{\mathcal{Z}}}$$\scriptstyle{\mu_{\mathcal{N}}}$$\scriptstyle{i_{\ast}^{fake}}$
Proof.
Observe that the contraction $\pi:\mathcal{N}\to\mathcal{Z}$ composed with $i^{fake}:\mathcal{Z}\hookrightarrow\mathcal{N}$ is homotopic to $id:\mathcal{N}\to\mathcal{N}$, thus induces a homotopy $\pi^{\ast}(i^{fake})^{\ast}\cong id$ on $\mathbb{C}(\mathcal{N})$. By [10] Theorem 3 in Section 3, we know that $i^{fake}_{\ast}\pi_{\ast}=id$ on $K_{0}^{fake}(\mathcal{N})$. On the other hand, it follows from $\pi i^{fake}=id_{\mathcal{Z}}$ that $\pi_{\ast}i^{fake}_{\ast}=id$ on $K_{0}^{fake}(\mathcal{Z})$. So $\pi_{\ast}$ and $i^{fake}_{\ast}$ are inverse to each other, and the corollary follows. ∎
Lemma 3.7.
Let $j:\mathcal{N}\to\mathcal{Y}$ be the inclusion of the closure of an open set.
  1. (1)
    $j^{!}j_{\ast}$ is the identity map $K_{0}^{fake}(\mathcal{N})\to K_{0}^{fake}(\mathcal{N})$.
  2. (2)
    If $\mathcal{Y}$ deformation retracts to $\mathcal{N}$, then $j^{!}$ and $j_{\ast}$ are inverse to each other.
Proof.
Part (1) follows from the definition. If furthermore $\mathcal{Y}$ deformation retracts to $\mathcal{N}$, then $j_{\ast}$ is an isomorphism by [10] Theorem 3 in Section 3. Thus $j_{\ast}=(j^{!})^{-1}$ by part (1). ∎
Proof of $i^{fake,!}$ being independent of the choice of $\mathcal{N}$.
Different tubular neighborhood structures on the same set $\mathcal{N}$ define the same map $i^{fake,!}$. The map $i^{fake,!}$ defined by $\mathcal{N}^{\prime}\subseteq\mathcal{N}$ and that defined by $\mathcal{N}$ coincide, by the second part of the previous lemma. Next, for two general tubular neighborhoods given, we can always find a tubular neighborhood contained in both. ∎
Lemma 3.8.
The following diagrams commutes for each open embedding $\mathcal{N}\hookrightarrow\mathcal{Y}$.
${K^{0,fake}(B^{\ast}\mathcal{N},\partial B^{\ast}\mathcal{N})}$${K^{0,fake}(B^{\ast}\mathcal{Y},\partial B^{\ast}\mathcal{Y})}$${K^{fake}_{0}(\mathcal{N})}$${K^{fake}_{0}(\mathcal{Y})}$$\scriptstyle{j_{\ast}}$$\scriptstyle{\mu_{\mathcal{N}}}$$\scriptstyle{\mu_{\mathcal{Y}}}$$\scriptstyle{j_{\ast}}$
${K^{0,fake}(B^{\ast}\mathcal{N},\partial B^{\ast}\mathcal{N})}$${K^{0,fake}(B^{\ast}\mathcal{Y},\partial B^{\ast}\mathcal{Y})}$${K^{fake}_{0}(\mathcal{N})}$${K^{fake}_{0}(\mathcal{Y})}$$\scriptstyle{j_{\ast}}$$\scriptstyle{\mu_{\mathcal{N}}}$$\scriptstyle{\mu_{\mathcal{Y}}}$$\scriptstyle{j^{!}}$
Proof.
The commutativity of the first diagram follows Corollary 2.2.
The second diagram’s commutativity follows that of the first diagram and the fact that $j!j_{a}st$ is the identity map. ∎
The well-definedness part of Proposition 3.3 follows from Lemma 3.5, Corollary 3.1, and Lemma 3.8.
3.4. Push-forward from inertia orbifold to point
Consider sequences $(g,g_{1}G_{1},\cdots,g_{n}G_{n})$ of cosets in $G$, where $G_{i+1}$ is the group generated by the previous term $g_{i}G_{i}$, and $G_{1}$ the group generated by $g$. Let $\mathcal{G}$ be such sequences with $g_{i}\notin G_{i}$. Given a sequence $\mathbf{g}=(g,g_{1},\cdots,g_{n})$, we can construct a sequence of such cosets by setting $G_{i}$ the group generated by $g,g_{1},\cdots,g_{i}$. The sequence of elements $\mathbf{g}$ defines an element in $\mathcal{G}$, if and only if each $g_{i}$ is not in the group generated by previous elements.
Given $\mathbf{g}$, let $\mathcal{Y}^{\mathbf{g}}$ be the $(n+1)$-multi-sector corresponding to $\mathbf{g}$, and $i_{\mathbf{g}}:\mathcal{Y}^{\mathbf{g}}\to\mathcal{Y}$. Let $n=|\mathbf{g}|$ and $m(\mathbf{g})$ be the order of isotropy of $\mathcal{Y}^{\mathbf{g}}$. Let $\mathcal{N}_{g_{i}}$ be the normal bundle of $\mathcal{Y}^{(g,g_{1},\cdots,g_{i})}$ in $\mathcal{Y}^{(g=g_{0},g_{1},\cdots,g_{i-1})}$, and $\mathcal{N}_{g}$ in particular be the normal bundle of $\mathcal{Y}^{(g)}$ in $\mathcal{Y}$. Define $e^{-1}_{\mathbf{g}}=(\prod_{i=0}^{n}str_{g_{i}}\wedge^{\ast}(\mathcal{N}^{\ast}_{g_{i}}\otimes\mathbb{C}))^{-1}$. These objects only depend on the element in $\mathcal{G}$ represented by $\mathbf{g}$, and are independent of the specific representative $\mathbf{g}$. We define the fake push-forward $p_{\ast}^{fake}$ of an element $(H,\phi,T)$ in the fake $K$-homology as the Fredholm index of $T_{+}:H_{+}\to H_{-}$, where $H_{\pm}$ are the positive and negative part of $H$ under the $\mathbb{Z}_{2}$ grading, and $T_{+}=T|_{H_{+}}$.
Definition 3.1.
The orbifold ‘push-forward to point’ $p_{\ast}(\varphi)$ of $\{\varphi^{(g)}=(H^{(g)},\phi^{(g)},T^{(g)})\}\in K_{0}(\mathcal{Y})$ is
$p_{\ast}(\varphi)=p_{\ast}^{fake}\left(\sum_{\mathbf{g}\in\mathcal{G}}\frac{(-1)^{|\mathbf{g}|+\dim\mathcal{Y}^{(g)}}}{m(\mathbf{g})}(e_{\mathbf{g}}^{-1}i_{\mathbf{g}}^{!}\varphi^{(g)})\right).$
Proposition 3.4.
The following diagram commutes, where $\mathbb{Z}\to\mathbb{C}$ is the natural inclusion, and $\operatorname{ind}(\cdots)_{+}$ maps $u$ to the index of $u_{+}$.
${{K^{0}(B^{\ast}\mathcal{Y},\partial B^{\ast}\mathcal{Y})}}$${K_{0}(\mathcal{Y})}$${\mathbb{Z}}$${\mathbb{C}}$$\scriptstyle{\mu}$$\scriptstyle{\operatorname{ind}(\cdots)_{+}}$$\scriptstyle{p_{\ast}}$
Proof.
Restricting or extending a symbol outside its support preserves the analytic index by the excision property (Theorem 3.7 in [3]). So taking into account Proposition 2.1, we only need to prove the proposition when $\mathcal{Y}=\mathcal{W}$ is closed. Given $u\in K^{0}(B^{\ast}\mathcal{Y},\partial B^{\ast}\mathcal{Y})$, let $u_{\mathbf{g}}=e_{\mathbf{g}}^{-1}i_{\mathbf{g}}^{\ast}(str_{g}u)$. With the assumption that $\mathcal{Y}$ is closed, we compute $\operatorname{ind}u_{+}$ by Kawasaki’s index theorem:
$\operatorname{ind}u_{+}=\sum_{(g)}(-1)^{\dim\mathcal{Y}^{(g)}}\int_{[T\mathcal{Y}^{(g)}]}ch(u_{g})_{+}\cdot td(T\mathcal{Y}^{(g)}\otimes\mathbb{C}).$
Note that supertraces act as identity on $u_{g}$.
Each term on the right-hand side is (up to a sign $(-1)^{\dim\mathcal{Y}^{(g)}}$) the principle term (that corresponds to the non-twisted strata $g_{1}\in\left<g\right>$, the group generated by $g$) of Kawasaki’s index formula for $u_{g}$ over $\mathcal{Y}^{(g)}$.
$\operatorname{ind}[u_{g}]_{+}=\sum_{(g_{1})}(-1)^{\dim\mathcal{Y}^{(g)(g_{1})}}\int_{[T\mathcal{Y}^{(g)(g_{1})}]}ch\left(\frac{i_{g_{1}}^{\ast}u_{g}}{str_{g_{1}}\wedge^{\ast}\mathcal{N}^{\ast}_{g_{1}}\otimes\mathbb{C}}\right)_{+}\cdot td(T\mathcal{Y}^{(g)(g_{1})}\otimes\mathbb{C}).$
Note that all terms with $\mathcal{Y}^{(g)(g_{1})}=\mathcal{Y}^{(g)}$ in the sum are the same. We move terms corresponding to $g_{1}\notin\left<g\right>$ to the left and divide both sides by $m(g)$ to get
$\begin{array}[]{ll}&\displaystyle\int_{[T\mathcal{Y}^{(g)(g_{1})}]}ch\left(\dfrac{i_{g_{1}}^{\ast}u_{g}}{str_{g_{1}}\wedge^{\ast}\mathcal{N}^{\ast}_{g_{1}}\otimes\mathbb{C}}\right)_{+}\cdot td(T\mathcal{Y}^{(g)(g_{1})}\otimes\mathbb{C})\\ =\dfrac{1}{m(g)}\operatorname{ind}[u_{g}]_{+}-&\displaystyle\sum_{(g,g_{1})\in\mathcal{G}}(-1)^{\dim\mathcal{Y}^{(g)(g_{1})}}\displaystyle\int_{[T\mathcal{Y}^{(g)(g_{1})}]}ch\left(\dfrac{i_{g_{1}}^{\ast}u_{g}}{str_{g_{1}}\wedge^{\ast}\mathcal{N}^{\ast}_{g_{1}}\otimes\mathbb{C}}\right)_{+}\cdot td(T\mathcal{Y}^{(g)(g_{1})}\otimes\mathbb{C}).\end{array}$
This expresses the desired integration on the sector $\mathcal{Y}^{(g)}$ in terms of the index of $u_{g}$ and integrations of the same form on $2$-sectors. The same result with $(\mathcal{Y}^{(\mathbf{g})},\mathcal{Y}^{(\mathbf{g})(g_{n+1})},u_{\mathbf{g}})$ in the place of $(\mathcal{Y}^{(g)},\mathcal{Y}^{(g)(g_{1})},u_{g})$ expresses relevant integrations on $(n+1)$-sectors in terms of $\operatorname{ind}[u_{\mathbf{g}}]_{+}$ and integrations on $(n+2)$-sectors. Since points in the compact orbifold $\mathcal{Y}$ have finite isotropy, there is a uniform bound $N$ on the order of the isotropy groups at points in $\mathcal{Y}$, and $n$-sectors $\mathcal{Y}^{\mathbf{g}}$ become empty when $n>N$. Thus, we can write
$ind([u]_{+})=\sum_{\mathbf{g}\in\mathcal{G}}\frac{(-1)^{|\mathbf{g}|+\dim\mathcal{Y}^{(g)}}}{m(\mathbf{g})}ind([u_{\mathbf{g}}]_{+})$
as a finite sum.
Next, we have
$(i_{\mathbf{g}})_{!}(u_{\mathbf{g}})|_{\mathcal{N}_{\mathbf{g}}}=str_{g}u|_{\mathcal{N}_{\mathbf{g}}}.$
This likewise follows from an induction on $|\mathbf{g}|$, where the base case and the induction step are both
$(i_{g})_{!}u_{(g)}|_{\mathcal{N}_{g}}=str_{g}u|_{\mathcal{N}_{g}}$
for appropriate $\mathcal{N},u$. Indeed, the latter is equivalent to
$\pi_{g}^{\ast}i_{g}^{\ast}\left(str_{g}\frac{u}{\wedge^{\ast}(\mathcal{N}_{g}^{\ast}\otimes\mathbb{C})}\right)\cdot str_{g}\wedge^{\ast}(\mathcal{N}_{g}^{\ast}\otimes\mathbb{C})=str_{g}u,$
which is
$\left(\pi_{g}^{\ast}i_{g}^{\ast}\bigg{(}str_{g}\frac{u}{\wedge^{\ast}(\mathcal{N}_{g}^{\ast}\otimes\mathbb{C})}\bigg{)}-\bigg{(}str_{g}\frac{u}{\wedge^{\ast}(\mathcal{N}_{g}^{\ast}\otimes\mathbb{C})}\bigg{)}\right)\cdot str_{g}\wedge^{\ast}(\mathcal{N}_{g}^{\ast}\otimes\mathbb{C})=0.$
The first term vanishes on $\mathcal{Y}^{g}$, and the second term vanishes on $\mathcal{N}^{g}\setminus\mathcal{Y}^{g}$.
Finally, applying part (1) of Proposition 3.3, we know that
$\begin{array}[]{ll}ind([u_{\mathbf{g}}]_{+})&=p_{\ast}^{fake}(\mu(u_{\mathbf{g}}))\\ &=p_{\ast}^{fake}i_{\mathbf{g}}^{!}\mu(i_{\mathbf{g}})_{!}(u_{\mathbf{g}})\\ &=p_{\ast}^{fake}(i_{\mathbf{g}}^{!}(\mu(u))^{(\mathbf{g})}),\end{array}$
and the proposition follows. ∎
3.5. Compatibility
Given an element $\alpha\in K^{0}(\mathcal{Y})$ and an element $T\in K_{0}(\mathcal{Y})$, we define their pairing
$T(\alpha)=p_{\ast}(\alpha\cap T).$
The following proposition follows and Propositions 3.2 and 3.4.
Proposition 3.5.
For $\beta\in K^{0}(B^{\ast}\mathcal{Y},\partial B^{\ast}\mathcal{Y})$ and $\alpha\in K^{0}(\mathcal{Y},\partial\mathcal{Y})$, we have
$(\mu(\beta))(\alpha)=p_{\ast}(\beta\boxtimes\alpha).$
4. K-theoretic virtual fundamental cycle
Definition 4.1.
A stably almost complex global Kuranishi chart consists of a quadruple $A=(G,Y,E,s)$, where $G$ is a compact Lie group, $Y$ a locally $C^{1}$ $G$-manifold with almost free $G$-action, such that $T^{\ast}_{G}Y$ is stably complex, $E$ a locally $C^{1}$ $G$-equivariant stably complex vector bundle over $Y$, and $s$ an equivariant section $Y\to E$.
In [2], Abouzaid, Mclean and Smith constructed $stably$ almost complex global Kuranishi charts on the compactified moduli space of stable maps ¯$\bar{\mathcal{M}}_{g.n}(X,d)$ to general symplectic targets $X$. We shall make the following modifications and use orbifold Kuranishi charts.
Definition 4.2.
An almost complex global orbifold Kuranishi chart consists of a triple $\mathcal{A}=(\mathcal{Y,E},\sigma)$, where $\mathcal{Y}$ a smooth, compact orbifold with boundary, $\mathcal{E}$ a complex smooth orbi-bundle over $\mathcal{Y}$, and $\sigma$ a continuous section $\mathcal{Y}\to\mathcal{E}$ that vanishes nowhere on $\partial\mathcal{Y}$.
Remark that almost complex global orbifold Kuranishi charts exist. Since the real dimensions of $Y$ and fibers of $E$ have the same parity, we can always add a same $\mathbb{R}^{k}$ to them (and identity map $id^{k}:\mathbb{R}^{k}\to\mathbb{R}^{k}$ to the section $s$) to get complex $Y,E$’s. An almost complex global Kuranishi chart $(G,Y,E,s)$ gives an almost complex orbifold Kuranishi chart $(\mathcal{Y,E},\sigma)$ by quotient out the $G$ action and shrinking to a compact neighborhood of $\sigma^{-1}(0)$. $C^{1}$ structures induce smooth structures, so we can also modify $Y,E$ to smooth orbifolds and smooth bundles. Note that $s$ is not smooth in general.
By [2] Proposition 4.62 and Lemma 3.4 (see also [1]), global Kuranishi charts $A=(G,Y,E,s)$ for the same moduli space of stable maps of a certain type are connected by a sequence of operations of the following types, or their inverses.
  1. (1)
    Free quotients $A\to A/H$, with $H$ a normal subgroup of $G$. This replaces $(G,Y,E,s)$ by $(G/H,Y/H,E/H,s)$, where $E/H$ is viewed as a $G/H$-bundle on $Y/H$.
  2. (2)
    The induced map $A\to A\times_{G}G^{\prime}$ associated to an embedding $G\to G^{\prime}$. This replaces $(G,Y,E,s)$ by $(G^{\prime},Y\times_{G}G^{\prime},\pi^{\ast}E,\pi^{\ast}s)$, where $\pi:Y\times_{G}G^{\prime}\to Y$ is the projection induced by a slicing $G^{\prime}\to G$.
  3. (3)
    The stabilization by a $G$-equivariant vector bundle $\pi:W\to Y$, replaces $(G,Y,E,s)$ by $(G,W,\pi^{\ast}E\oplus\pi^{\ast}W,\pi^{\ast}s\oplus\Delta)$, where $\Delta$ is the tautological ‘diagonal’ section of $W\to\pi^{\ast}W$.
  4. (4)
    An open embedding $A^{\prime}\to A$ of a neighborhood of $s^{-1}(0)$. This replaces $(G,Y^{\prime},E|_{Y^{\prime}},s|_{Y^{\prime}})$ by $(G,Y,E,s)$ for $Y^{\prime}\subseteq Y$ open containing $s^{-1}(0)$.
In turn, the orbifold Kuranishi charts are related by a sequence of the following operation and its inverse.
[…]
We define K-theoretic virtual fundamental cycle $\mathcal{O}^{vir}_{\mathcal{Y,E},\sigma}\in K_{0}(\mathcal{Y})$ of an almost complex global orbifold Kuranishi chart $(\mathcal{Y,E},\sigma)$ as follows. Let $\pi:T^{\ast}\mathcal{Y}\to\mathcal{Y}$ be the projection that forgets the cotangent vector. Let $\mathcal{S}=\wedge^{\ast}_{\mathbb{C}}T^{\ast}\mathcal{Y},\theta=\wedge^{\ast}_{\mathbb{C}}\mathcal{E}$ and $\mathcal{O}=\mathcal{S}\otimes\theta$. Remark that $\mathcal{S},\theta$ are $\mathbb{Z}_{2}$-graded bundles and so is $\mathcal{E}$. We add Hermitian structures to $\mathcal{Y,E}$, and let
$\begin{array}[]{ll}p&=\lambda_{\xi}\otimes 1+\varepsilon 1\otimes\pi^{\ast}\lambda_{\sigma}:\pi^{\ast}\mathcal{O}\to\pi^{\ast}\mathcal{O},\\ \mathcal{O}^{vir}_{\mathcal{Y,E},\sigma}&=\mu(p)\end{array}.$
Recall that $\varepsilon$ is the operator on $\mathbb{Z}_{2}$ graded objects acting as $(-1)^{a}$ on the direct summand of degree $a$. Also recall that $\lambda$ is as defined in the introduction, namely given a section $\tau:\mathcal{X}\to\mathcal{F}$ of a Hermitian bundle, we define $\lambda_{\tau}:\wedge^{\ast}\mathcal{F}\to\wedge^{\ast}\mathcal{F}$ by $\tau(x)\wedge\cdot+i_{\tau(x)}(\cdot)$ over points $x\in\mathcal{X}$, where $i_{\tau(x)}$ is the contraction with $\tau(x)$ via the Hermitian metric on $\mathcal{F}$. Form $\lambda_{\xi}:\pi^{\ast}\mathcal{S}\to\pi^{\ast}\mathcal{S}$ and $\lambda_{\sigma}:\mathcal{E}\to\mathcal{E}$ corresponding to the diagonal map $\xi:T^{\ast}\mathcal{Y}\to\pi^{\ast}T^{\ast}\mathcal{Y}$ and $\sigma:\mathcal{Y}\to\mathcal{E}$. Observe that $\lambda_{\xi}$ (resp. $\lambda_{\sigma}$) are of degree $1$, so $p$ is also of degree $1$. Observe that $\lambda_{\xi},\lambda_{\sigma}$ are self-dual and squares to $|\xi|^{2}\cdot 1,|\sigma(x)|^{2}\cdot 1$, so $p^{\ast}p=(|\xi|^{2}+|\sigma(x)|^{2})\cdot 1$. Hence $p$ is invertible outside the compact set =$\mathcal{M}=\{|\sigma(x)|=0,|\xi|=0\}$ and therefore represents an element in $K^{0}(T^{\ast}\mathcal{Y})$.
It follows from Proposition 3.3 that in terms of the K-theoretic virtual fundamental cycle, the operation (*) corresponds to the push forward by $i_{\ast}$ (defined in Section 2.3) and then pull back by $j^{!}$ (defined in Section 3.3).
Theorem 1.
If $\sigma$ is transversal to the zero section 00, then $\mathcal{O}^{vir}_{\mathcal{Y,E},\sigma}=i_{\ast}[\sigma^{-1}(0)]$.
Proof.
Note that the space $\mathcal{M}=\sigma^{-1}(0)$ is stably almost complex. We can define its K-theoretic fundamental cycle by Bott periodicity theorem $K_{0}(\Sigma^{k}\mathcal{M})=K_{0}(\mathcal{M})$. This identification is exactly our $i^{!}$ for $i:\mathcal{M}\to B(\mathcal{M}\times\mathbb{R}^{2k})$. Next, by part (2) of Proposition 3.3, the push-forwards $i_{\ast}$ and the wrong-way pull-back $i^{!}$ are intertwined by the choose-an-operator map, thus proving the theorem. ∎
[…]
Theorem 2.
$\mathcal{O}^{vir}_{\sigma},\mathcal{V}\right>=\operatorname{ind}[\mathcal{E}\otimes\pi^{\ast}\mathcal{V}]_{+}.$
Corollary 4.1.
The invariants are independent of the global Kuranishi chart chosen.